• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Armstrong

    Log entry time 16:10:28 on October 23, 2009

    Entry number 296324

    keyword=Shift Summary

    Shift Summary   Day Shift   October 23 2009   (HAPPEX-III)

    D. Armstrong (SL)
    K. Rider (TO)
    S. Frullani (3rd)

    8:00 Status at start of shift:
    - Beam off, waiting for Target cooldown to begin

    8:03 MCC calls - they will send 5 µA of CW down our line to work on the FFB (Fast FeedBack) system. Dave Meekins arrives, and he will start the target cool down as soon as the FFB system work is done.

    10:30 MCC has fixed Fast Feedback, and have tested it at 5 µA and at 100 µA. Compton rates are bad at high current, but decide not to have MCC try to track that down, go next to target cooldown. Call for controlled access. Dave Meekins goes into Hall. Alex C. and Rupesh S. also make access to work on Compton and HAPPEX detector issues respectively.

    15:45 Target is cold and filled. Many thanks to Dave M! Everyone is now out of the hall...

    16:00 Shift ends as we are trying to get all the HRS magnets back to operating point...

    No data runs taken....